Approximately how long will it take for your money to triple if it is invested at 4.5% compounded continously?
Answer by stanbon(57387)

Approximately how long will it take for your money to triple if it is invested at 4.5% compounded continously?
---------
A(t) = P*e^(rt)
---
3P = P*e^(0.045t)
---
e^(0.045t) = 3
---
Take the natural log of both sides to to get:
0.045t = ln(3)
t = ln(3)/0.045
---
t = 24.41 years
======================
Cheers,
Stan H.
----------------